# Synergistic eigenanalysis of covariance and Hessian matrices for enhanced binary classification on health datasets

### v2

@@ A.3.2 Variance Preservation Theorem for Projection onto a Vector
A.3.3 Variance Ratio Preservation Theorem for Projection onto a Vector
Consider an arbitrary subset of points located on the first axis, denoted by $X=\{x_{1},x_{2},\ldots,x_{n}\}$, with variances $\sigma_{X_{1}}^{2}$ and $\sigma_{X_{2}}^{2}$ for subsets $X_{1}$ and $X_{2}$, respectively. These points are projected onto the vector $\mathbf{v}=(v_{1},v_{2})$ using the projection operation defined earlier.
The projected subsets of points on $\mathbf{v}$ are denoted by $Y_{1}=\{y_{1},y_{2},\ldots,y_{n}\}$ and $Y_{2}=\{z_{1},z_{2},\ldots,z_{n}\}$, where each $y_{i}$ and $z_{i}$ represents the projection of $x_{i}$ onto $\mathbf{v}$. The variances of $Y_{1}$ and $Y_{2}$, denoted as $\sigma_{Y_{1}}^{2}$ and $\sigma_{Y_{2}}^{2}$ respectively, measure the spread of the projected points around their respective means.
@@ A.3.3 Variance Ratio Preservation Theorem for Projection onto a Vector
The Variance Ratio Preservation Theorem states that the ratio of variances between the projected subsets $Y_{2}$ and $Y_{1}$ is equal to the ratio of variances between the original subsets $X_{2}$ and $X_{1}$.
Mathematically, this can be expressed as:
$\frac{\sigma_{Y_{2}}^{2}}{\sigma_{Y_{1}}^{2}}=\frac{\sigma_{X_{2}}^{2}}{\sigma_{X_{1}}^{2}}$
Proof:
By the Variance Preservation Theorem for Projection onto a Vector, we know that the variance of the projected subset $Y$ can be expressed as:
$\sigma_{Y}^{2}=v_{1}^{2}\cdot\sigma_{X}^{2}$
Applying this theorem to $Y_{1}$ and $X_{1}$, we have:
$\sigma_{Y_{1}}^{2}=v_{1}^{2}\cdot\sigma_{X_{1}}^{2}$
Similarly, applying the theorem to $Y_{2}$ and $X_{2}$, we have:
$\sigma_{Y_{2}}^{2}=v_{1}^{2}\cdot\sigma_{X_{2}}^{2}$
Now, we can take the ratio of these variances:
$\frac{\sigma_{Y_{2}}^{2}}{\sigma_{Y_{1}}^{2}}=\frac{v_{1}^{2}\cdot\sigma_{X_{2}}^{2}}{v_{1}^{2}\cdot\sigma_{X_{1}}^{2}}=\frac{\sigma_{X_{2}}^{2}}{\sigma_{X_{1}}^{2}}$
Therefore, we have shown that the ratio of variances between the projected subsets $Y_{2}$ and $Y_{1}$ is equal to the ratio of variances between the original subsets $X_{2}$ and $X_{1}$, as required.
@@
To extend Theorem 1 to multidimensional data, we consider a dataset where each sample is represented by a $D$-dimensional column vector $\vec{\theta}=\left[\theta_{1},\theta_{2},\cdots,\theta_{D}\right]^{T}$. The dataset of $2n$ samples is divided into two subsets: $C_{1}=\{\vec{x}_{i}\in\mathbf{R}^{D}\mid i=1,2,\ldots,n\}$ and $C_{2}=\{\vec{y}_{j}\in\mathbf{R}^{D}\mid j=1,2,\ldots,n\}$, each containing $n$ samples. The data in $C_{1}$ and $C_{2}$ follow the same underlying distribution, centered around their respective means, $\vec{m}_{1}$ and $\vec{m}_{2}$, resulting in the combined data from $C_{1}$ and $C_{2}$ being centered around an overall mean of $\vec{m}$. The covariances for $C_{1}$ and $C_{2}$ are equal, noted as $Cov_{1}=Cov_{2}=Cov_{w}$, and the combined data has an overall covariance of $Cov$. The individual distributions for the vectors $\vec{X}$ and $\vec{Y}$ are isotropic about their respective means, meaning their covariances are proportional to the unit matrix in D dimensions, i.e., $Cov_{w}=\sigma_{w}^{2}I_{D}$. The distance between the class means, $d=\left\|\left(\vec{m}_{1}-\vec{m}_{2}\right)\right\|$, indicates the separation between $\vec{m}_{1}$ and $\vec{m}_{2}$.
Define the sample means
$\vec{m}=\frac{1}{2n}\sum_{k=1}^{2n}\vec{\theta}_{k}=\frac{n}{2n}\vec{m}_{1}+\frac{n}{2n}\vec{m}_{2}=\frac{1}{2}\left(\vec{m}_{1}+\vec{m}_{2}\right)$ (B.1)
@@
$\vec{m}_{1}=\frac{1}{n}\sum_{i=1}^{n}\vec{x}_{i}\text{ and }\vec{m}_{2}=\frac{1}{n}\sum_{j=1}^{n}\vec{y}_{j}$ (B.2)
@@
The biased sample covariance matrix of the combined training set is then given by
$Cov=\frac{1}{2n}\sum_{k=1}^{2n}\left(\vec{\theta}_{k}-\vec{m}\right)\left(\vec{\theta}_{k}-\vec{m}\right)^{T}=\frac{1}{2n}\sum_{k=1}^{2n}\vec{\theta}_{k}\vec{\theta}_{k}^{T}-\vec{m}\vec{m}^{T}$ (B.3)
Expanding this yields
$Cov=\frac{1}{2n}\left[\sum_{i=1}^{n}\vec{x}_{i}\vec{x}_{i}^{T}+\sum_{j=1}^{n}\vec{y}_{j}\vec{y}_{j}^{T}\right]-\vec{m}\vec{m}^{T}$ (B.4)
Now,
$\sum_{i=1}^{n}\vec{x}_{i}\vec{x}_{i}^{T}=\sum_{i=1}^{n}\vec{x}_{i}\vec{x}_{i}^{T}-\left(n\right)\vec{m}_{1}\vec{m}_{1}^{T}+\left(n\right)\vec{m}_{1}\vec{m}_{1}^{T}=n\left(Cov_{1}+\vec{m}_{1}\vec{m}_{1}^{T}\right)$ (B.5)
and similarly
$\sum_{j=1}^{n}\vec{y}_{j}\vec{y}_{j}^{T}=n\left(Cov_{2}+\vec{m}_{2}\vec{m}_{2}^{T}\right)$ (B.7)
@@
where $C_{1}$ and $C_{2}$ are the (biased) sample covariances of the class 1 and class 2 subsets respectively. Substituting (B.1), (B.5), and (B.7) in (B.4), after some algebra, leads to
$Cov=\frac{n}{2n}Cov_{1}+\frac{n}{2n}Cov_{2}+\frac{n^{2}}{\left(2n\right)^{2}}\left(\vec{m}_{1}-\vec{m}_{2}\right)\left(\vec{m}_{1}-\vec{m}_{2}\right)^{T}=Cov_{w}+\frac{1}{4}\left(\vec{m}_{1}-\vec{m}_{2}\right)\left(\vec{m}_{1}-\vec{m}_{2}\right)^{T}$ (B.8)
It reduces to the result of Theorem 1 for $\mathrm{D}=1$. To see this, take the trace of (B.8) to show that
$\operatorname{tr}(Cov)=\operatorname{tr}\left(Cov_{w}\right)+\frac{1}{4}\operatorname{tr}\left(\left(\vec{m}_{1}-\vec{m}_{2}\right)\left(\vec{m}_{1}-\vec{m}_{2}\right)^{T}\right)$ (B.9)
But $\operatorname{tr}\left(\left(\vec{m}_{1}-\vec{m}_{2}\right)\left(\vec{m}_{1}-\vec{m}_{2}\right)^{T}\right)=\left\|\vec{m}_{1}-\vec{m}_{2}\right\|^{2}$ and the trace of a diagonalizable matrix is equal to the sum of its eigenvalues, in this case the sum of the variances of the corresponding training set along the directions of the eigenvectors. So, if $D=1$ then
$\sigma^{2}=\frac{1}{2}\sigma_{1}^{2}+\frac{1}{2}\sigma_{2}^{2}+\frac{1}{4}\left(m_{1}-m_{2}\right)^{2}=\sigma_{\omega}^{2}+\frac{1}{4}\left(m_{1}-m_{2}\right)^{2}$ (B.10)
@@
which is equivalent to the result of Theorem 1. Theorem 1, however, is supposed to hold for the projection of the data onto an arbitrary 1D subspace of the full D-dimensional data space, and to imply that maximizing $\sigma^{2}$ by choosing the direction of this subspace appropriately will also be the direction which maximizes the difference between the projected sample means. Fortunately, since $Cov_{w}=\sigma^{2}I_{D}$, we have
$Cov=\sigma_{w}^{2}I_{D}+\frac{1}{4}\left(\vec{m}_{1}-\vec{m}_{2}\right)\left(\vec{m}_{1}-\vec{m}_{2}\right)^{T}$ (B.11)
Multiplying both sides with $\left(\vec{m}_{1}-\vec{m}_{2}\right)$ yields
$Cov\left(\vec{m}_{1}-\vec{m}_{2}\right)=\left(\sigma_{w}^{2}+\frac{1}{4}d^{2}\right)\left(\vec{m}_{1}-\vec{m}_{2}\right)$ (B.12)
@@
Please note that equation (B.12) forms an eigenequation where $Cov$ is the matrix of interest, the term $\left(\sigma_{w}^{2}+\frac{1}{4}d^{2}\right)$ represents the eigenvalue, and $\left(\vec{m}{1}-\vec{m}{2}\right)$ is the eigenvector. This implies that the difference between the mean vectors $\left(\vec{m}{1}-\vec{m}{2}\right)$ is indeed an eigenvector of the full set’s covariance matrix $Cov$. This alignment confirms that the projection maximizing the variance in the D-dimensional space will also maximize the between-class mean distance, ensuring the validity of Theorem 1 in higher dimensions.

### v3

@@ A.3 Variance ratio preservation property for projection onto a vector
Theorem 4 (Variance ratio preservation upon projection onto a vector).
Let $X=\{x_{1},x_{2},\ldots,x_{n}\}$ be a set of 1D data points, with two subsets $X_{1}=\{x_{1},\ldots,x_{m}\}$ and $X_{2}=\{x_{m+1},\ldots,x_{n}\}$, having variances $\sigma_{X_{1}}^{2}$ and $\sigma_{X_{2}}^{2}$, respectively. Consider projecting $X$ onto a unit vector $\mathbf{v}=(v_{1},v_{2})$ in a 2-dimensional space, resulting in a set of projected points $Y=\{y_{1},y_{2},\ldots,y_{n}\}$, with corresponding projected subsets $Y_{1}$ and $Y_{2}$.
@@
Then the ratio of variances between the two projected subsets $Y_{1}$ and $Y_{2}$ is equal to the ratio of the variances between their original subsets $X_{1}$ and $X_{2}$:
$\frac{\sigma_{Y_{2}}^{2}}{\sigma_{Y_{1}}^{2}}=\frac{\sigma_{X_{2}}^{2}}{\sigma_{X_{1}}^{2}}.$
Proof.
For a point $(x_{i},0)$ on the first axis, its projection onto $\mathbf{v}=(v_{1},v_{2})$ is given by $y_{i}=x_{i}\cdot v_{1}$. Consequently, the distance between two projected points $y_{i}$ and $y_{j}$ is $|y_{i}-y_{j}|=|x_{i}\cdot v_{1}-x_{j}\cdot v_{1}|=|v_{1}|\cdot|x_{i}-x_{j}|$. Thus, the distances between pairs of points in the original space and their projections are scaled uniformly by $|v_{1}|$.
Next, consider the variance of the projected subset $Y_{1}$, denoted as $\sigma_{Y_{1}}^{2}$. We can compute it using the formula $\sigma_{Y_{1}}^{2}=\frac{1}{m-1}\sum_{i=1}^{m}(y_{i}-\mu_{Y_{1}})^{2}$, where $\mu_{Y_{1}}$ is the mean of the subset $Y_{1}$. Substituting $y_{i}=x_{i}\cdot v_{1}$, and using the fact that $(y_{i}-\mu_{Y_{1}})=v_{1}\cdot(x_{i}-\mu_{X_{1}})$, we obtain $\sigma_{Y_{1}}^{2}=v_{1}^{2}\cdot\frac{1}{m-1}\sum_{i=1}^{m}(x_{i}-\mu_{X_{1}})^{2}=v_{1}^{2}\cdot\sigma_{X_{1}}^{2}$.
Similarly, the variance for the second subset $Y_{2}$ is given by $\sigma_{Y_{2}}^{2}=v_{1}^{2}\cdot\sigma_{X_{2}}^{2}$.
Taking the ratio of the variances of the projected subsets $Y_{1}$ and $Y_{2}$:
$\frac{\sigma_{Y_{2}}^{2}}{\sigma_{Y_{1}}^{2}}=\frac{v_{1}^{2}\cdot\sigma_{X_{2}}^{2}}{v_{1}^{2}\cdot\sigma_{X_{1}}^{2}}=\frac{\sigma_{X_{2}}^{2}}{\sigma_{X_{1}}^{2}}.$
This establishes that the ratio of variances is preserved in the projection, as required. ∎
@@
To extend Theorem 1 to multidimensional data, we consider a dataset where each sample is represented by a $D$-dimensional column vector $\vec{\theta}=\left[\theta_{1},\theta_{2},\cdots,\theta_{D}\right]^{T}$. The dataset of $2n$ samples is divided into two subsets: $C_{1}=\{\vec{x}_{i}\in\mathbf{R}^{D}\mid i=1,2,\ldots,n\}$ and $C_{2}=\{\vec{y}_{j}\in\mathbf{R}^{D}\mid j=1,2,\ldots,n\}$, each containing $n$ samples. The data in $C_{1}$ and $C_{2}$ follow the same underlying distribution, centered around their respective means, $\vec{\mu}_{1}$ and $\vec{\mu}_{2}$, resulting in the combined data from $C_{1}$ and $C_{2}$ being centered around an overall mean of $\vec{\mu}$. The covariances for $C_{1}$ and $C_{2}$ are denoted as $S_{1}$ and $S_{2}$, respectively, and the combined data has an overall covariance of $S$. The individual distributions for the vectors $\vec{X}$ and $\vec{Y}$ are isotropic about their respective means, meaning their covariances are proportional to the unit matrix in $D$ dimensions, i.e., $S_{1}=\sigma_{1}^{2}I_{D}$ and $S_{2}=\sigma_{2}^{2}I_{D}$. The distance between the class means, $d=\left\|\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)\right\|$, indicates the separation between $\vec{\mu}_{1}$ and $\vec{\mu}_{2}$.
Define the sample means:
$\vec{\mu}=\frac{1}{2n}\sum_{k=1}^{2n}\vec{\theta}_{k}=\frac{n}{2n}\vec{\mu}_{1}+\frac{n}{2n}\vec{\mu}_{2}=\frac{1}{2}\left(\vec{\mu}_{1}+\vec{\mu}_{2}\right)$ (8)
@@
$\vec{\mu}_{1}=\frac{1}{n}\sum_{i=1}^{n}\vec{x}_{i}\text{ and }\vec{\mu}_{2}=\frac{1}{n}\sum_{j=1}^{n}\vec{y}_{j}$ (9)
@@
The biased sample covariance matrix of the combined training set is then given by:
$S=\frac{1}{2n}\sum_{k=1}^{2n}\left(\vec{\theta}_{k}-\vec{\mu}\right)\left(\vec{\theta}_{k}-\vec{\mu}\right)^{T}=\frac{1}{2n}\sum_{k=1}^{2n}\vec{\theta}_{k}\vec{\theta}_{k}^{T}-\vec{\mu}\vec{\mu}^{T}$ (10)
Expanding this yields:
$S=\frac{1}{2n}\left[\sum_{i=1}^{n}\vec{x}_{i}\vec{x}_{i}^{T}+\sum_{j=1}^{n}\vec{y}_{j}\vec{y}_{j}^{T}\right]-\vec{\mu}\vec{\mu}^{T}$ (11)
Now:
$\sum_{i=1}^{n}\vec{x}_{i}\vec{x}_{i}^{T}=\sum_{i=1}^{n}\vec{x}_{i}\vec{x}_{i}^{T}-\left(n\right)\vec{\mu}_{1}\vec{\mu}_{1}^{T}+\left(n\right)\vec{\mu}_{1}\vec{\mu}_{1}^{T}=n\left(S_{1}+\vec{\mu}_{1}\vec{\mu}_{1}^{T}\right)$ (12)
and similarly:
$\sum_{j=1}^{n}\vec{y}_{j}\vec{y}_{j}^{T}=n\left(S_{2}+\vec{\mu}_{2}\vec{\mu}_{2}^{T}\right)$ (13)
@@
where $S_{1}$ and $S_{2}$ are the (biased) sample covariances of the class 1 and class 2 subsets respectively. Substituting (8), (12), and (13) in (11), after some algebra, leads to:
$S=\frac{n}{2n}S_{1}+\frac{n}{2n}S_{2}+\frac{n^{2}}{\left(2n\right)^{2}}\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)^{T}=\frac{1}{2}S_{1}+\frac{1}{2}S_{2}+\frac{1}{4}\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)^{T}$ (14)
@@
Theorem 1 is supposed to hold for the projection of the data onto an arbitrary 1D subspace of the full D-dimensional data space, and to imply that maximizing $\sigma^{2}$ by choosing the direction of this subspace appropriately will also be the direction which maximizes the difference between the projected sample means. Fortunately, since $S_{1}=\sigma_{1}^{2}I_{D}$ and $S_{2}=\sigma_{2}^{2}I_{D}$, we have:
$S=\frac{1}{2}\sigma_{1}^{2}I_{D}+\frac{1}{2}\sigma_{2}^{2}I_{D}+\frac{1}{4}\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)^{T}$ (15)
Multiplying both sides with $\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)$ yields:
$S\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)=\left(\frac{1}{2}\sigma_{1}^{2}+\frac{1}{2}\sigma_{2}^{2}+\frac{1}{4}d^{2}\right)\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)$ (16)
@@
Please note that equation (16) forms an eigenequation where $S$ is the matrix of interest, the term $\left(\frac{1}{2}\sigma_{1}^{2}+\frac{1}{2}\sigma_{2}^{2}+\frac{1}{4}d^{2}\right)$ represents the eigenvalue, and $\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)$ is the eigenvector. This implies that the difference between the mean vectors $\left(\vec{\mu}_{1}-\vec{\mu}_{2}\right)$ is indeed an eigenvector of the full set’s covariance matrix $S$. This alignment confirms that the projection maximizing the variance in the D-dimensional space will also maximize the between-class mean distance, ensuring the validity of Theorem 1 in higher dimensions.
@@ Appendix C Additional dataset results
Appendix D Within-class covariance matrices for unnormalized WBCD dataset
Figure 14: Absolute covariance matrices for unnormalized WBCD dataset by class. Panel (a) presents the absolute within-class covariance matrices for unnormalized WBCD dataset in decimal scale. It reveals that the variances are dominated by a couple of components, each on the order of $10^{5}$, along with their significant covariance. These notable variances and covariance overshadow those of the other components, some of which are on the order of $10^{-5}$. To enhance the visualization of variances and covariance structures, panel (b) utilizes a log scale. However, even with this adjustment, the covariance matrices reveal clusters without displaying a consistent pattern of uniform diagonal elements or isomorphism, clearly demonstrating that the lack of normalization has contributed to this irregular structure.